OpenStudy (anonymous):

I need step by step instructions on how to simplify 6 1/2

OpenStudy (anonymous):

\[\Large 6\cdot \frac12 =\frac62=3\]

OpenStudy (diyadiya):

\[6 \frac{1}{2}= 6+ \frac{1}{2} \]Taking LCM of 1 and 2\[\frac{6}{1}+ \frac{1}{2}= \frac{6 \times 2}{ 1 \times 2}+ \frac{1}{2}\]\[\frac{12}{2}+ \frac{1}{2}\]\[\frac{12+1}{2}= \frac{13}{2}\]

OpenStudy (anonymous):

aufff.. sorry... I thought it was multiply :S

OpenStudy (anonymous):

@NinaCarter is it ...\[\Large 6\frac12 \quad ,\quad 6\cdot \frac12 \quad ,\quad 6+\frac12\] ?

OpenStudy (anonymous):

Thank you now how do I simplify the square root of 100?

OpenStudy (anonymous):

@Kreshnik it is the first one (six and a half)

OpenStudy (diyadiya):

\[\sqrt{100}=\sqrt{10 \times 10}=\sqrt{10^2}=10\] Remember \[\sqrt{x^2}=x\]

OpenStudy (anonymous):

if it is the first one then you'll go like thiss... \[\Large 6\frac12 =\frac{2\cdot 6+1}{2}=\frac{8+1}{2}=\frac92\]

OpenStudy (anonymous):

wrong ... I made a mistake 2*6=12 then... 12+1=13 finally 13/2 LOL

OpenStudy (anonymous):

@diyadia and Kreshnik, this was very helpful. Have a great day :)

OpenStudy (diyadiya):

glad to help :)